# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ F.3 Universality of massive emission
We now show that the relations to the right of (F.25) and (F.26) can in fact be directly inferred from the NR expression of the emission amplitudes of the relevant processes. In the following, we consider only the case of distinguishable particle scattering with long-range mediation [^18] The short-range mediation case follows the same arguments, with an $r$-expansion instead. It additionally includes the contact diagrams shown in Fig.1(c), whose emission amplitudes reduce to the same analytic expressions as the one from external-leg emissions., where the $m_{\phi,V}$-corrections arises exclusively from the four external leg emission diagrams, Fig.1(a). We restrict ourselves to the two diagrams in which the DM species 1 emits the boson, since the remaining diagrams follow by the simple substitutions $a_{1}\leftrightarrow a_{2}$, $m_{1}\leftrightarrow m_{2}$, $E_{1},E_{3}\to E_{2},E_{4}$, and $p_{1},p_{3}\to p_{2},p_{4}$ applied to the amplitude pieces defined below.
For a given scenario $k$, the emission amplitude contributions (or, for $V$-emission, its 0-component) associated to an emission from the external legs of the DM species 1, reads
@@
Here, $E_{3}=m_{1}+{|\mathbf{p}_{f}|^{2}}/{2m_{1}}+\mathcal{O}(v_{i}^{3})$ and $E_{1}=m_{1}+{|\mathbf{p}_{i}|^{2}}/{2m_{1}}+\mathcal{O}(v_{i}^{4})$. For $S_{1}S_{2}\to S_{1}S_{2}V$, the 0-component of the currents are found by applying the substitution $\delta_{s_{1},s_{3}}\delta_{s_{2},s_{4}}\to 1$ in (F.30).
@@ F.4 Scattering of identical particles
$\left\langle|\mathcal{M_{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=$ $\frac{64y^{6}}{15\omega^{2}}|\vec{k}|^{2}|\vec{l}|^{2}\left\{\frac{\left(1+6(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+\kappa^{2}\left(2+7(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+\frac{\kappa^{4}}{4}\left(4+19(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)}{|\vec{k}|^{4}}+(\mathbf{k}\leftrightarrow\mathbf{l})\right$. (F.34)
$-\frac{1}{|\vec{k}|^{2}|\vec{l}|^{2}}\Bigg{[}\left(1+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)+\kappa^{2}\left(2+3(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)$
@@
$\left\langle|\mathcal{M_{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=$ $\left(\frac{A}{m_{S}}\right)^{6}\frac{|\vec{k}|^{2}|\vec{l}|^{2}}{15\omega^{2}}\left\{\frac{\left(1+11(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+2\kappa^{2}\left(1+6(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+\frac{\kappa^{4}}{4}\left(4+19(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)}{|\vec{k}|^{4}}+(\mathbf{k}\leftrightarrow\mathbf{l})\right$. (F.35)
$\left.+2\frac{1}{{|\vec{k}|^{2}|\vec{l}|^{2}}}\Bigg{[}\left(1+9(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)+2\kappa^{2}\left(1+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)\right$.

### v2

@@ F.3 Universality of massive emission
We now show that the relations to the right of (F.25) and (F.26) can in fact be directly inferred from the NR expression of the emission amplitudes of the relevant processes. In the following, we consider only the case of distinguishable particle scattering with long-range mediation [^19] The short-range mediation case follows the same arguments, with an $r$-expansion instead. It additionally includes the contact diagrams shown in Fig.1(c), whose emission amplitudes reduce to the same analytic expressions as the one from external-leg emissions., where the $m_{\phi,V}$-corrections arise exclusively from the four external leg emission diagrams, Fig.1(a). We restrict ourselves to the two diagrams in which the DM species 1 emits the boson, since the remaining diagrams follow by the simple substitutions $a_{1}\leftrightarrow a_{2}$, $m_{1}\leftrightarrow m_{2}$, $E_{1},E_{3}\to E_{2},E_{4}$, and $p_{1},p_{3}\to p_{2},p_{4}$ applied to the amplitude pieces defined below.
For a given scenario $k$, the emission amplitude contributions (or, for $V$-emission, its 0-component) associated with an emission from the external legs of the DM species 1, read
@@
Here, $E_{3}=m_{1}+{|\mathbf{p}_{f}|^{2}}/{2m_{1}}+\mathcal{O}(v_{i}^{3})$ and $E_{1}=m_{1}+{|\mathbf{p}_{i}|^{2}}/{2m_{1}}+\mathcal{O}(v_{i}^{4})$. For $S_{1}S_{2}\to S_{1}S_{2}V$, the 0-component of the currents are found by applying the substitution $\delta_{s_{1},s_{3}}\delta_{s_{2},s_{4}}\to 1$ in (F.30); with this replacement, the contributions from vertex emission are effectively accounted for.
@@ F.4 Scattering of identical particles
$\left\langle|\mathcal{M_{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=$ $\frac{64y^{6}}{15\omega^{2}}|\vec{k}|^{2}|\vec{l}|^{2}\left\{\frac{\left(1+6(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)-\kappa^{2}\left(2+7(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+\frac{\kappa^{4}}{4}\left(4+19(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)}{|\vec{k}|^{4}}+(\mathbf{k}\leftrightarrow\mathbf{l})\right$. (F.34)
$-\frac{1}{|\vec{k}|^{2}|\vec{l}|^{2}}\Bigg[\left(1+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)-\kappa^{2}\left(2+3(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)$
@@
$\left\langle|\mathcal{M_{Q}}|^{2}\right\rangle_{\mathbf{\bm{\hat{}}{\mathbf{q}}}}=$ $\left(\frac{A}{m_{S}}\right)^{6}\frac{|\vec{k}|^{2}|\vec{l}|^{2}}{15\omega^{2}}\left\{\frac{\left(1+11(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)-2\kappa^{2}\left(1+6(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)+\frac{\kappa^{4}}{4}\left(4+19(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}\right)}{|\vec{k}|^{4}}+(\mathbf{k}\leftrightarrow\mathbf{l})\right$. (F.35)
$\left.+2\frac{1}{{|\vec{k}|^{2}|\vec{l}|^{2}}}\Bigg[\left(1+9(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)-2\kappa^{2}\left(1+4(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{2}+2(\bm{\hat{}}{\mathbf{k}}\cdot\bm{\hat{}}{\mathbf{l}})^{4}\right)\right$.